OpenStudy (adrianna.gongora):

How do I convert y = a(x –7)^2 + 28 into standard form?

OpenStudy (photon336):

Step #1 you need to expand that (x-7)^2 and foil it you know how to do that? \[a(x-7)(x-7)+28 \]

OpenStudy (mathmale):

Please expand (x-7)^2, and then mult. each term of the result by the coefficient a.

OpenStudy (mathmale):

Thus: a(x-7)^2=a(x^2 - ???? ) +28

OpenStudy (adrianna.gongora):

\[(x ^{2} +7x+ 7x+ 14) + 28\]

OpenStudy (mathmale):

Wouldn't each "7x" actually be "-7x?" Where did that +14 come from? (Hint: What's 7^2)?)

OpenStudy (adrianna.gongora):

I foiled a(x−7)(x−7)+28

OpenStudy (adrianna.gongora):

@mathmale

OpenStudy (adrianna.gongora):

(x^2-7x-7x+49)+28

OpenStudy (mathmale):

Much better. Combining terms: x^2 - 14x + 49 +28 = ??

OpenStudy (adrianna.gongora):

14x^3+77

OpenStudy (mathmale):

We must keep the powers of x separate. x^2 - 14x cannot be combined. The correct result would be x^2 - 14x + 77.

OpenStudy (adrianna.gongora):

I thought so but I wasn't sure so I just combined them.

OpenStudy (adrianna.gongora):

So is x^2 - 14x + 77 standard form of y=a(x-7)^2+28?

OpenStudy (mathmale):

Originally you had y = a(x-7)^2+28. We can't just forget about the 'a." So now we have to figure out how to include the 'a' in your final equation. You could begin with your intermediate result, x^2 - 14x + 49 (no 28) and mult. each of these three terms by 'a.' This would end with 49a, which can't be combined with 28 except by writing "49a+28."

OpenStudy (adrianna.gongora):

so 49a + 28 is the final answer ?

OpenStudy (adrianna.gongora):

@Satellite73

Directrix (directrix):

>>so 49a + 28 is the final answer ? No.

OpenStudy (adrianna.gongora):

I'm confused so what do I do do I times (x^2 - 14x + 49) (49a +28)

Directrix (directrix):

y = a * ( x^2 - 49x + 49 ) + 28 To this point is correct.

Directrix (directrix):

Distribute (multiply) the "a" times this: ( x^2 - 49x + 49 ). Do not multiply a times the 28. It is not inside the parentheses with the other terms.

Directrix (directrix):

>>I times (x^2 - 14x + 49) (49a +28) No

OpenStudy (adrianna.gongora):

y = a * ( x^2 - 49x + 49 ) + 28 so this would be the correct answer or would I need to multiply this?

Directrix (directrix):

What is a times 49 =